If i had 16000 invested into two accounts one yielding 4% and the other 10% and received a total of 1,060 at the end of the year. How much did i invest into each account.
Answer by mananth(16946)

Investment I 4.00% per annum ---x
Investment II 10.00% per annum ---y

x + y= 16000 ------------------------1
4.00% x + 10.00% y= = $1,060.00
Multiply by 100
4 x + 10 y= = $106,000.00

Multiply (1) by -4
we get

-4 x -4 y= = -64000.00

Add this to (2)

0 x 6 y= = $42,000.00

divide by 6

y = $7,000.00
Balance $9,000.00
CHECK
$9,000.00 --------- 4.00% ------- $360.00
$7,000.00 ------- 10.00% ------- $700.00
Total -------------- $1,060.00
m.ananth@hotmail.ca